ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 82, 1993

Aschbacher, M. Simple cennectivity of p-group complexes  1

Cohen, A. M. and D. B. Wales Embeddings of the group L(2,13) in groups of Lie type E_{6}

45

Farkas, H. M. and I. Kra Automorphic forms for subgroups of the modular group

87

Farkas, H. M. and Y. Kopeliovich New theta constant identities 133
Feit, W. The K-admissibility of 2A_{6} and 2A_{7}  141

Fried, M. D., R. Guralnick and J. Saxl Schur covers and Carlitz’s conjecture

157

Gorenstein, D. and R. Lyons On Aschbacher’s local C(G; T) theorem

227

Guralnick, R. M. and G. R. Robinson On extensions of the Baer– Suzuki Theorem

281

Guralnick, R. See M. D. Fried

Hartley, B. and A. E. Zalesskiĭ On simple periodic linear groups— dense subgroups, permutation representations, and induced modules

299

Herzog, M., P. Longobardi and M. Maj On a combinatorial problem in group theory

329

Ivanov, A. A. and S. V. Shpectorov The last flag-transitive P -geometry

341

Kopeliovich, Y. See H. M. Farkas

Kra, I. See H. M. Farkas

Longobardi, P. See M. Herzog

Lubotzky, A., A. Mann and D. Segal Finitely generated groups of polynomial subgroup growth

363

Lyons, R. See D. Gorenstein

Maj, M. See M. Herzog

Mann, A. Some properties of polynomial subgroup growth groups  373

Mann, A. See A. Lubotzky

Robinson, G. R. See R. M. Guralnick

Saxl, J. See M. D. Fried Segal, D. See A. Lubotzky

Segev, Y. Group actions on finite acyclic simplicial complexes  381
Shalev, A. Automorphisms of finite groups of bounded rank  395

Shpectorov, S. V. See A. A. Ivanov Völklein, H. Braid group action via GL_{n}(q) and U_{n}(q), and Galois

realizations

405

Wales, D. B. See A. M. Cohen Zalesskiĭ, A. E. See B. Hartley